OpenStudy (anonymous):

Q(t) = 4+12sin((π/30)t) t ε [0,30] Find the global max'm

OpenStudy (campbell_st):

well \[Q'(t) = 12 \times \pi/30 \cos(\pi/30 t)\] \[Q'(t) = 3\pi/5 \cos(\pi/30 t)\] Q'(t) = 0 and solve for t